The sum of two numbers is 65 and their difference is 5. What are the two numbers? Write answer as an ordered pair: (x, y)

Answer by sheldonbbtrocks(115)

This is a basic systems of equations problem.
1. Write the problem in 2 equations:
x+y=65
x-y=5
Assuming x is the larger number and y is the smaller number
2. Cancel either the x's or y's multiply an equation if necessary:
We don't need to multiply any equations, because the y's already cancel to 0. We can now add the x's and the numbers on the outside:
2x=70
Solve for x
x=35
3. Substitute x=35 into any of the original equations (let's do the first one):
35+y=65
Solve for y:
y=65-35
y=30
4: Check
35+30=65
35-30=5
Both equations are correct!
Therefore, your ordered pair is [(35,30)]